[Hallb-engineering] [New Logentry] RC Report for Saturday, December 4, 2021

schmidta at jlab.org schmidta at jlab.org
Sat Dec 4 12:50:02 EST 2021

Logentry Text:
Last 24 hours:
- Production running on carbon at 90 nA continues to be mostly smooth
- Sergey continues to work on improving the stability of the DAQ. We have switched back to the PROD66 configuration, and plan to use that moving forward.
- The epicsweb accelerator status had a proxy error at 2AM, and this has affected our displays in the counting house.
- We had a little down time last night due to a 3rd pass separator trip, but we're back up and running by 7:20 this AM. 

Coming up:
- We plan to keep taking production data on carbon until tomorrow AM
- Starting around 8AM, we'll initiate the switch from C to Sn. This will entail re-centering the beam on the new target, and doing a luminosity scan to establish production current.
- Monday, we are scheduled for a full day of down time to install the Ca-48 target. 

- I have been working to restore full-size displays in the counting house for the accelerator status and the beamline status. However, my efforts have been in vain (see attachment) If anyone knows how to get these displays to be BIG, your help would be greatly appreciated. 

That's it for this report. I will make a similar report tomorrow after the target change. As always, if I am missing anything, don't be shy about giving me a call. 
Wishing you a good weekend, and pleasant beam,


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3956278
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallb-engineering/attachments/20211204/cb06b11d/attachment.html>

More information about the Hallb-engineering mailing list